# Spin-chain multichannel Kondo model via image impurity boundary condition

### v1

@@ Abstract
One of the signature observables for the electronic multichannel Kondo model is the impurity entropy, which was claimed to be found in $J_{1}$-$J_{2}$ Heisenberg chains with the open boundary condition (OBC) and periodic boundary condition (PBC), for the one-channel and two-channel cases respectively. However, it is not clear how to generalize OBC and PBC in Heisenberg chains to find the multichannel Kondo impurity entropy with more than two channels. In this paper, we demonstrate that the correct boundary condition for realizing multichannel Kondo physics in Heisenberg chains is the image impurity boundary condition (IIBC) which yields the expected impurity entropy, $\ln[(\sqrt{5}+1)/2]$ for the three-channel case and $\ln\sqrt{3}$ for the four-channel case. Moreover, the IIBC reduces to OBC for the one-channel case and to PBC for the two-channel case. With IIBC, the finite-size scaling of the impurity entropy and the total impurity spin both match the finite-temperature corrections in the electronic multichannel Kondo model. Additionally, we show that the XXZ anisotropy lowers the impurity entropy by a power law of the effective Luttinger liquid parameter.
@@ I Introduction
The electronic Kondo effect occurs at low temperatures when conduction electrons screen a magnetic impurity leading to a many-body entangled state [1, 2]. This model was later generalized to consider multiple free-fermion channels overscreening the impurity [3], and this generalization has proved to be a useful model to illustrate exotic many-body phenomena [4, 5, 6], especially in mesoscopic devices [7, 8, 9, 10, 11]. A striking example is the zero-temperature impurity entropy $S_{\text{imp}}=\ln g$ where $g$ plays the role of a “ground-state degeneracy” which is generally not an integer [12, 5, 6]. For a Kondo model with $M$ channels, it has been shown through the Bethe Ansatz method [13, 14] and by boundary conformal field theory methods [12, 5, 15, 6, 16] that:
@@
Recently, the multichannel Kondo model has also attracted the attention of the topological quantum computation community since $g(M)$ matches the quantum dimension of various non-Abelian anyons [17, 18], offering the tempting interpretation that an emergent anyon appears at the overscreened impurity. The case of $M=3$ is of particular interest since $g(M=3)=\varphi\approx 1.618$, the golden ratio, hints at the existence of Fibonacci anyons [17, 18, 16, 19, 20], exotic quasiparticles that may provide a path to universal topological quantum computation [21]. However, the precise connection to anyons remains unclear and, so far, an explicit construction is only known for the Majorana case $M=2$ [22, 23], motivating further investigation of $g(M>2)$.
@@
While experiments towards observing impurity entropy and verifying Eq. (1) are progressing [24, 25, 26], this formula has been numerically verified for up to $M=3$ channels via the chiral numerical renormalization group (NRG) method [27, 28]. Theoretically, conformal field theory methods, specifically non-abelian bosonization, have shown that the free fermions of the electronic $M$-channel Kondo model can be bosonized using the “charge” [$U(1)_{2M}$], “spin” [$\text{SU}(2)_{M}$] and “channel” [$\text{SU}(M)_{2}$] currents in the corresponding Kac-Moody algebras [15, 5, 16]. The Kondo interaction is only coupled with the spin sector, suggesting that the fermionic background is not necessary and that Kondo physics can also be studied in systems such as spin chains [29, 30, 31] or quantum circuits [32].
@@
Furthermore, it has been widely known that the low-energy effective theory of the spin-chain Heisenberg (XXX) model is described by the $\text{SU}(2)_{1}$ Kac-Moody algebra [33]. Indeed, the one-channel and two-channel Kondo impurity entropies, that is Eq. (1) at $M=1,2$, were observed in the open boundary and periodic boundary XXX spin-chain models, respectively [29, 34]. These spin-chain models are of interest because they directly demonstrate that a fermionic background is not necessary to realize Kondo physics and thus that the essence of the Kondo effect may indeed be bosonic. Additionally, the use of a spin-chain model enables us to take advantage of the DMRG algorithm [35] which has seen great success in quantum simulation [36, 37, 38]. However, as we illustrate below, it is not obvious how to extend the spin chain formulation to $M>2$ channels.
Even the realization of the two-channel model using an $\text{SU}(2)_{1}$ Heisenberg chain is subtle. One considers two spin chains, one of size $L+1$ for the first channel, and one of size $L$ for the second, in which spins are uniformly coupled with a strength $J$. The use of channels of two different lengths was originally to make the total number of sites $1+N_{L}+N_{R}$ even [34]. On one end of the chains, one uses a coupling constant $J_{K}$ to attach the two chains together with an impurity spin at $x=0$. On the other end, the two spin chains are coupled to each other with coupling $J$, thereby “healing” the chains together [39]. Thus, when $J_{K}=J$, the two-channel spin-chain system reduces to a uniform, periodic chain of length $22L+$ 2. This property allows one to extract the impurity entropy at the Kondo intermediate coupling fixed point.
@@
Figure 1: (a) The geometry of the spin-chain $M$-channel Kondo model (next-nearest neighbor coupling $J_{2}$ not shown). An impurity (left) and an image impurity (right) are both connected to the $M$ bulk channels but with different couplings $J_{K}$ and $J_{K}^{\prime}$. The nearest-neighbor couplings in the bulk channel are always $J_{1}$. The total number of spins is $M\times L+2$. The cut position is labeled by $x=1,\dots,L-1$. (b-c) The entanglement entropy as a function of $x$ for $J_{K}=J_{K}^{\prime}=J_{1}$, $J_{2}=J_{2}^{c}=0.2412J$ obtained from DMRG simulation. The uniform part $S_{u}$ is close to the prediction $M\times S^{\text{OBC}}+\ln g(M)$ where from Eq. (1) $g(3)=\varphi$ is the golden ratio and $g(4)=\sqrt{3}$. In (b) $L=100$ and DMRG maximum bond dimension $\chi_{\text{max}}=4000$ while in (c) $L=40$ and $\chi_{\text{max}}=5000$.
@@
where $c=1$ is the central charge, $s_{1}$ is a non-universal factor, and $g_{\text{Cardy}}=2^{-1/4}$ is the non-integer ground-state degeneracy associated with a Heisenberg chain [12]. Notably, for the two-channel model ($M=2$), when $J_{K}=J_{1}$, $S^{2\text{CK}}_{u}(x)$ is given by the Cardy formula for a periodic chain of length $22L+$ 2 with a cut position $12x+$ 1 [see Fig. 1(a)],
@@
The value of $g=\sqrt{2}$ coincides with the prediction of Eq. 1 for $M=2$ and matches the quantum dimension of Majorana fermions. That said, it is important to note that $g=\sqrt{2}$ is not unique to the two-channel spin-chain Kondo model [41, 42, 43].
@@
Additionally, we also consider the effects of exchange anisotropy on the spin-chain multichannel Kondo models. One-channel [50] and two-channel Kondo physics [41] have also been observed in XX Heisenberg chains where the next-nearest neighbor coupling $J_{2}$ is not needed, unlike the isotropic case [29, 34]. Notably, the absence of the next-nearest neighbor interaction makes numerical simulation much easier, motivating potential investigation into the viability of XX chains for simulating Kondo physics [45]. Furthermore, one-channel Kondo physics has also been demonstrated for a general anisotropic 1D antiferromagnetic Heisenberg chain [51]. Following these investigations, we introduce anisotropy into our multichannel Kondo model, parametrized by $\Delta$ with $\Delta=1$ being the isotropic model and $\Delta=0$ being the XX model. We numerically observe that $g\sim d^{(M-1)/2}$ where $d=1-\arccos(\Delta)/\pi$ (see Fig. 3) and the effective Luttinger liquid parameter of the XXZ chain is $1/(2d)$ [52], motivating further investigations.
II The Multichannel Spin-Chain Kondo Model
$H=$ $\ H_{\text{imp}}+H_{\text{image}}+\sum_{k=1}^{M}H^{k}_{\text{bulk}}\,,$ (6)
$H^{k}_{\text{bulk}}\!=\!$ $\ J_{1}\sum_{j=1}^{L-1}{\bf\it\mathrm{S}}_{k,j}\cdot{\bf\it\mathrm{S}}_{k,j+1}+J_{2}\sum_{j=1}^{L-2}{\bf\it\mathrm{S}}_{k,j}\cdot{\bf\it\mathrm{S}}_{k,j+2}\,,$ (7)
$H_{\text{imp}}\!=\!$ $\ J_{K}\!\sum_{k=1}^{M}[{\bf\it\mathrm{S}}_{0}\!\cdot\!\quantity(J_{1}{\bf\it\mathrm{S}}_{k,1}\!+\!J_{2}{\bf\it\mathrm{S}}_{k,2})\!+\!J_{2}\sum_{k^{\prime}>k}{\bf\it\mathrm{S}}_{k,1}\!\cdot\!{\bf\it\mathrm{S}}_{k^{\prime},1}],$ (8)
where the spin-1/2 operator ${\bf\it\mathrm{S}}_{k,j}$ is ${\bf\it\mathrm{S}}=\hbar(\sigma^{x},\sigma^{y},\sigma^{z})/2$ for the $j^{\text{th}}$ spin in $k^{\text{th}}$ channel. $H_{\text{imp}}$ and $H_{\text{image}}$ implement interaction with and across the impurity and image impurity respectively where $H_{\text{image}}$ is defined analogously to $H_{\text{imp}}$ but with $J_{K}\to J_{K}^{\prime}$. The impurity is coupled with a coupling $J_{K}$. The couplings between spins are illustrated in Fig. 1(a). For simplicity, we work in units $J_{1}=\hbar=1$.
The next-nearest neighbor couplings $J_{2}$ are introduced to suppress a marginally irrelevant bulk interaction which contributes to the finite size effects [53, 54, 55, 56, 34]. These finite size effects are minimized when the marginal coupling constant in the bulk vanishes which has been shown to occur at the critical point $J^{c}_{2}\approx 0.2412$ [55] between the gapless ($J_{2}<J^{c}_{2}$) and dimerized ($J_{2}>J^{c}_{2}$) phases. Note further that we fix the chain length $L$ to be even as it has been shown that an odd chain length leads to quantitatively different entanglement entropy profiles which do not match Kondo physics [57, 58, 59].
III Impurity Entropy
@@
In this work, we use the entanglement entropy as a probe for the $M$-channel Kondo model. We use DMRG (TenPy [60]) to numerically find the ground state of Eq. (6) and extract the entanglement entropy $S(x)$. The finite system size $L$ induces oscillations in $S(x)$ which allow us to decompose $S(x)$ into a uniform and alternating term [56], $S(x)=S_{u}(x)+(-1)^{x}S_{a}(x)$. For the one-channel Kondo model, the bosonization of the spin-chain Kondo Hamiltonian reveals that this alternating term corresponds to a staggered spin density operator which has no counterpart in the electronic Kondo model [29]. Similar oscillations also occur for the two-channel model when $J_{K}\neq 1$ [34]. Thus, in this work, we focus on the uniform part $S_{u}(x)$ for our analysis of the entanglement structure. Given the entropy profile $S(x)$, we can extract $S_{u}(x)$ via a cubic fit, developed initially for one-channel models in [57].
We then calculate the impurity entropy $S_{\text{imp}}(x)$ as in Eqs. (2)-(3). The case of $J_{K}=1$ approximately corresponds to the Kondo fixed point [29] wherein the contribution to the impurity entropy is a consequence of the spin-chain geometry itself. In the thermodynamic limit $L\to\infty$, we expect the contribution to the impurity entropy to be constant throughout the chain with small deviations at the boundaries [56, 59]. The constant impurity entropy is expressed as $\ln g$ where the effective ground state degeneracy $g$ is predicted to be given by Eq. 1 [12].
@@
We first present our numerical values of $g$ for the $M=3$ and $M=4$ models. We compute $g$ from $\exp\quantity(S_{\text{imp}})$ at $x=L/2$ which corresponds to the closest agreement with Eq. 1 as shown in Fig. 1(b-c). In computing $S_{\text{imp}}$, we also must compute $S^{\text{OBC}}$ which requires the determination of the non-universal constant $s_{1}$. Fortunately, $s_{1}$ can be easily extracted by comparing DMRG results with Eq. 4 for a periodic XXX chain. Based on this $s_{1}$ value, we find that, at the Kondo fixed point ($J_{K}=1$), $g\approx 1.623$ for $M=3$ and $g\approx 1.787$ for $M=4$. Both these values are very close to the predictions of Eq. 1, i.e., $g(M=3)=1.618$ and $g(M=4)=1.732$, even for the relatively small system sizes we used. The difference can be attributed to finite size effect as we will discuss below.
When $J_{K}\to 0$, we expect the screening of the impurity to decrease corresponding to an increase in the entanglement entropy [29, 34]. In this case, the entanglement entropy will no longer be uniform and will instead increase with cut position as the impurity is approached. General predictions for multichannel Kondo physics suggest that $S_{\text{imp}}(x\ll L)\to\ln 2$ as $J_{K}\to 0$ corresponding to the formation of a so-called impurity valence bond (IVB) between the impurity and the bulk [57, 61]. The additional many-body entanglement induced by this IVB then corresponds to the Kondo screening cloud which has a length scale set by the Kondo correlation length $\xi_{K}$. It has been predicted for the single-channel case that $\xi_{K}$ should scale as $e^{a/J_{K}}$ for $a>0$ [62, 2]. Moreover, the impurity entropy $S_{\text{imp}}$ is predicted to be a scaling function of $x/L$ and $x/\xi_{K}$ [57, 59]. The dependence of $\xi_{K}$ and scaling of $S_{\text{imp}}$ have also been observed for the two-channel model [34].
@@
We seek to further support our proposed multichannel model by showing that it reproduces a similar scaling of $S_{\text{imp}}$ and $\xi_{K}$ observed for the one- and two-channel models. We performed an analysis of $S_{\text{imp}}$ as $J_{K}$ was varied for different values of the system size $L$ to extract $\xi_{K}$ by fitting to the scaling form $S_{\text{imp}}(x/L,x/\xi_{K})$. The extracted $\xi_{K}$ values are then shown in the inset of Fig. 2(a) where our empirically measured $\xi_{K}$ indeed appears to scale as $e^{a/J_{K}}$. The resulting scaling law for $S_{\text{imp}}$ is shown in Fig. 2(a). In contrast to the Fermi-liquid power law $1/(x/\xi_{K})$ observed in the spin-chain 1CK model [57, 59, 29] and the non-Fermi liquid power law $\ln(x/\xi_{K})/(x/\xi_{K})$ for the spin-chain 2CK model [34], we find the non-Fermi liquid power law $1/(x/\xi_{K})^{4/5}$ in our spin-chain 3CK model, which is more dominant than the Fermi liquid $1/(x/\xi_{K})$ coming from each XXX chain. Both finite-size power laws match the finite-temperature correction to the multichannel impurity entropy [63, 64], which is related to the specific heat [16] when taking the temperature $T\to 1/(x/\xi_{K})$. Fig. 2(a) allows us to estimate the finite-size correction $\exp S_{\text{imp}}(x)-g(M)$. For $M=3$ and the parameters of Fig. 1(b), we can estimate this correction to be ∼$\sim 0.01$ for $x=L/2$. Similar estimate for $M=4$ (albeit using the above $\xi_{K}$) gives ∼$\sim 0.03$, of the same order of magnitude as $\exp S_{\text{imp}}-\sqrt{3}$ in Fig. 1(c). Similar to above finite-size scaling is found for the total impurity spin, discussed next.
@@
IV Total Impurity Spin

### v2

@@ Abstract
One of the signature observables for the electronic multichannel Kondo model is the impurity entropy, which was found in $J_{1}$-$J_{2}$ Heisenberg chains with the open boundary condition (OBC) and periodic boundary condition (PBC), for the one-channel and two-channel cases respectively. However, it is not clear how to generalize OBC and PBC in Heisenberg chains to find the multichannel Kondo impurity entropy with more than two channels. In this paper, we demonstrate that the correct boundary condition for realizing multichannel Kondo physics in Heisenberg chains is the image impurity boundary condition (IIBC) which preserves reflection symmetry and yields the expected impurity entropy, $\ln[(\sqrt{5}+1)/2]$ for the three-channel case and $\ln\sqrt{3}$ for the four-channel case. Moreover, the IIBC reduces to OBC for the one-channel case and to PBC for the two-channel case. With IIBC, the finite-size scaling of the impurity entropy and the total impurity spin match the finite-temperature corrections in the electronic multichannel Kondo model. Additionally, we show dependence of the impurity entropy, the total impurity spin, and their scaling behaviors on the XXZ anisotropy $\Delta$ (equivalently the Luttinger liquid parameter), revealing impurity physics in a multichannel Luttinger liquid.
@@ I Introduction
The electronic Kondo effect occurs at low temperatures when conduction electrons screen a magnetic impurity leading to a many-body entangled state [1, 2]. This model was later generalized to consider multiple free-fermion channels overscreening the impurity [3], and this generalization has proved to be a useful model to illustrate exotic many-body phenomena [4, 5, 6], especially in mesoscopic devices [7, 8, 9, 10, 11], many of which exhibit non-Fermi liquid behavior. A striking example is the zero-temperature impurity entropy $S_{\text{imp}}=\ln g$ where $g$ plays the role of a “ground-state degeneracy” which is generally not an integer [12, 5, 6]. For a Kondo model with $M$ channels, it has been shown through the Bethe Ansatz method [13, 14] and by boundary conformal field theory methods [12, 5, 15, 6, 16] that:
@@
Note that $g(M=1)=1$ gives zero entropy due to exact screening and $g(M=2)=\sqrt{2}$ (the quantum dimension of a free Majorana) gives $\ln(2)/2$ entropy due to overscreening [17].
@@
Recently, the multichannel Kondo model has also attracted the attention of the topological quantum computation community since $g(M)$ matches the quantum dimension of various non-Abelian anyons [18, 19], offering the tempting interpretation that an emergent anyon appears at the overscreened impurity. The case of $M=3$ is of particular interest since $g(M=3)=\varphi\approx 1.618$, the golden ratio, hints at the existence of Fibonacci anyons [18, 19, 16, 20, 21], exotic quasiparticles that may provide a path to universal topological quantum computation [22]. However, the precise connection to anyons remains unclear and, so far, an explicit construction is only known for the Majorana case $M=2$ [17, 23], motivating further investigation of $g(M>2)$ in multichannel Kondo models. Beyond the pure multichannel Kondo models, Ref. [24] explicitly finds the $\mathbb{Z}_{3}$ parafermion in the two-impurity Kondo model [11].
@@
While experiments towards observing impurity entropy and verifying Eq. (1) are progressing [25, 26, 27], this formula has been numerically verified for up to $M=4$ channels via various numerical renormalization group (NRG) methods [28, 29, 30, 31]. Theoretically, conformal field theory methods, specifically non-abelian bosonization, have shown that the free fermions of the electronic $M$-channel Kondo model can be bosonized using the “charge” $U(1)_{2M}$], “spin” $\text{SU}(2)_{M}$] and “channel” $\text{SU}(M)_{2}$] currents in the corresponding Kac-Moody algebras [15, 5, 16]. The Kondo interaction is only coupled with the spin sector, suggesting that the fermionic background is not necessary and that Kondo physics can also be studied in systems such as spin chains [32, 33, 34] or quantum circuits [35].
@@
Furthermore, it has been widely known that the low-energy effective theory of the spin-chain Heisenberg (XXX) model is described by the $\text{SU}(2)_{1}$ Kac-Moody algebra [36]. Indeed, the one-channel and two-channel Kondo impurity entropies, that is Eq. (1) at $M=1,2$, were observed in the open boundary and periodic boundary XXX spin-chain models, respectively [32, 37]. These spin-chain models are of interest because they directly demonstrate that a fermionic background is not necessary to realize Kondo physics. Additionally, the use of a spin-chain model enables us to take advantage of the DMRG algorithm [38], which has seen great success in quantum simulation [39, 40, 41]. Spin-chain models also complement fermionic models simulated by NRG. However, as we illustrate below, it is not obvious how to extend the spin chain formulation to $M>2$ channels. This extension is one of the key results in this letter and serves to suggest that the physics of the Kondo effects may be bosonic.
Besides the $\text{SU}(2)$-symmetric case, we also consider the effects of exchange anisotropy on the spin-chain multichannel Kondo models. One-channel [42] and two-channel Kondo physics [43] have also been observed in XX Heisenberg chains where the next-nearest neighbor coupling $J_{2}$ is not needed, unlike the isotropic case [32, 37]. Notably, the absence of the next-nearest neighbor interaction makes numerical simulation much easier, motivating potential investigation into the viability of XX chains for simulating Kondo physics [44]. Furthermore, one-channel Kondo physics has also been demonstrated for a general anisotropic 1D antiferromagnetic Heisenberg chain [45]. Following these investigations, we introduce anisotropy into our multichannel Kondo model, parametrized by $\Delta$ with $\Delta=1$ being the isotropic model and $\Delta=0$ being the XX model. We numerically observe that $g\sim d^{(M-1)/2}$ where $d=1-\arccos(\Delta)/\pi$ (see Fig. 3). Here, anisotropy is related to the effective Luttinger liquid parameter of the XXZ chain $K=1/(2d)$ [46], motivating further investigations on the entanglement in multichannel Luttinger liquid impurity models [47, 48, 49, 50, 51, 52, 53, 54].
II The Multichannel Spin-Chain Kondo Model
$H=\ H_{\text{imp}}+H_{\text{image}}+\sum_{k=1}^{M}H^{k}_{\text{bulk}}\,$, (2)
$H^{k}_{\text{bulk}}\!=\!\ J_{1}\sum_{j=1}^{L-1}{\bf\it\mathrm{S}}^{k}_{j}\cdot{\bf\it\mathrm{S}}^{k}_{j+1}+J_{2}\sum_{j=1}^{L-2}{\bf\it\mathrm{S}}^{k}_{j}\cdot{\bf\it\mathrm{S}}^{k}_{j+2}\,$, (3)
$H_{\text{imp}}\!=\!\ J_{K}\!\sum_{k=1}^{M}\Big[{\bf\it\mathrm{S}}_{0}\!\cdot\!\quantity({\bf\it\mathrm{S}}^{k}_{1}\!+\!\frac{J_{2}}{J_{1}}{\bf\it\mathrm{S}}^{k}_{2})\!+\!\frac{J_{2}}{J_{1}}\sum_{k^{\prime}>k}{\bf\it\mathrm{S}}^{k}_{1}\!\cdot\!{\bf\it\mathrm{S}}^{k^{\prime}}_{1}\Big],$ (4)
$H_{\text{image}}\!=\!\ J_{K}^{\prime}\!\sum_{k=1}^{M}\Big[{\bf\it\mathrm{S}}_{0}^{\prime}\!\cdot\!\quantity({\bf\it\mathrm{S}}^{k}_{L}\!+\!\frac{J_{2}}{J_{1}}{\bf\it\mathrm{S}}^{k}_{L-1})\!+\!\frac{J_{2}}{J_{1}}\sum_{k^{\prime}>k}{\bf\it\mathrm{S}}^{k}_{L}\!\cdot\!{\bf\it\mathrm{S}}^{k^{\prime}}_{L}\Big],$ (5)
where ${\bf\it\mathrm{S}}^{k}_{j}$ is the spin-1/2 operator ${\bf\it\mathrm{S}}=\hbar(\sigma^{x},\sigma^{y},\sigma^{z})/2$ for the $j^{\text{th}}$ spin in $k^{\text{th}}$ channel. $H_{\text{imp}}$ and $H_{\text{image}}$ implement interaction with and across the impurity and image impurity respectively. The nearest-neighbor couplings between spins are also illustrated in Fig. 1(a). For simplicity, we work in units $J_{1}=\hbar=1$. To implement anisotropy, we use the Hamiltonian Eq. 2 but replace ${\bf\it\mathrm{S}}_{i}\cdot{\bf\it\mathrm{S}}_{j}\to\mathrm{S}_{i}^{x}\mathrm{S}_{j}^{x}+\mathrm{S}_{i}^{y}\mathrm{S}_{j}^{y}+\Delta\mathrm{S}_{i}^{z}\mathrm{S}_{j}^{z}$. We consider only the critical (gapless) regime $1\!<\!\Delta\!\leq\!1$ [46]. The isotropic case $\Delta\!=\!1$ gives the multichannel spin-chain Kondo model where the Kondo impurity entropy matches Eq. (1). Note that the $\Delta=0$ case is a free-fermion model only for $M=1,2$, which further demonstrates that the generalization of Fig. 1(a) is nontrivial.
The next-nearest neighbor couplings $J_{2}$ are introduced to suppress a marginally irrelevant bulk interaction which contributes to finite size effects [55, 56, 57, 58, 37]. These finite size effects are minimized when the marginal coupling constant in the bulk vanishes, which has been shown to occur at the critical point $J^{c}_{2}\approx 0.2412$ [57] between the gapless ($J_{2}<J^{c}_{2}$) and dimerized ($J_{2}>J^{c}_{2}$) phases. Note further that we fix the chain length $L$ to be even as it has been shown that an odd chain length leads to quantitatively different entanglement entropy profiles which do not match Kondo physics [59, 60, 61].
III The Subtle $M=2$ Case
Even the realization of the two-channel model [see Fig. 1(a)] using an $\text{SU}(2)_{1}$ Heisenberg chain is subtle. Following Ref. [37], we consider a uniform, periodic chain of length $2L+2$ ($J_{K}=J^{\prime}_{K}=J_{1}$) to extract the impurity entropy at the Kondo intermediate coupling fixed point.
@@ III The Subtle $M=2$ Case
Figure 1: (a) The geometry of the spin-chain $M$-channel Kondo model (next-nearest neighbor coupling $J_{2}$ not shown). An impurity (left) and an image impurity (right) are both connected to the $M$ bulk channels but with different couplings $J_{K}$ and $J_{K}^{\prime}$. The nearest-neighbor couplings in the bulk channel are always $J_{1}$. The next nearest-neighbor couplings $J_{2}$ are omitted in the figure for simplicity, see Eqs. (3)-(5) for details. The total number of spins is $M\times L+2$. The cut position is labeled by $x=1,\dots,L-1$. The entanglement entropy as a function of $x$ for $J_{K}=J_{K}^{\prime}=J_{1}$, $J_{2}=J_{2}^{c}=0.2412J_{1}$ obtained from DMRG simulation for (b) $M=3$ and (c) $M=4$. The uniform part $S_{u}$ is close to the prediction $M\times S^{\text{OBC}}+\ln g(M)$ where from Eq. (1) $g(3)=\varphi$ is the golden ratio and $g(4)=\sqrt{3}$. In (b) $L=100$ and DMRG maximum bond dimension $\chi_{\text{max}}=4000$ while in (c) $L=40$ and $\chi_{\text{max}}=5000$.
@@
where $c=1$ is the central charge, $s_{1}$ is a non-universal factor, and $g_{\text{Cardy}}=2^{-1/4}$ is the non-integer ground-state degeneracy associated with a Heisenberg chain [12]. Notably, for the two-channel model ($M=2$), when $J_{K}=J_{1}$, $S^{2\text{CK}}_{u}(x)$ is given by the Cardy formula for a periodic chain of length $2L+2$ with a cut position $2x+1$ [see Fig. 1(a)],
@@
The value of $g=\sqrt{2}$ coincides with the prediction of Eq. 1 for $M=2$ and matches the quantum dimension of Majorana fermions. That said, it is important to note that $g=\sqrt{2}$ is not unique to the single-impurity two-channel spin-chain Kondo model [43, 63, 64], much less to the two-impurity Kondo model [65, 66].
@@
IV Numerical Impurity Entropy
@@
In this work, we use the entanglement entropy as a probe for the $M$-channel Kondo model. We use DMRG (TenPy [72, 73]) to numerically find the ground state of Eq. (2) and extract the entanglement entropy $S(x)$. The finite system size $L$ induces oscillations in $S(x)$ which allow us to decompose $S(x)$ into a uniform and alternating term [58], $S(x)=S_{u}(x)+(-1)^{x}S_{a}(x)$. In this work, we focus on the uniform part $S_{u}(x)$ for our analysis of the entanglement structure [32, 37]. Given the entropy profile $S(x)$, we can extract $S_{u}(x)$ as the average of two cubic fits for the upper blue dots (odd $x$) and lower blue dots (even $x$) [59], see Fig. 1(b-c). We then calculate the impurity entropy $S_{\text{imp}}(x)$ as in Eqs. (6)-(7). The case of $J_{K}=1$ approximately corresponds to the Kondo fixed point [32, 37] wherein the contribution to the impurity entropy is a consequence of the spin-chain geometry itself.
@@
We first present our numerical values of $g$ for the $M=3$ and $M=4$ models. We compute $g$ from $\exp\quantity(S_{\text{imp}})$ at $x=L/2$ which corresponds to the closest agreement with Eq. 1 as shown in Fig. 1(b-c). In computing $S_{\text{imp}}$, we also must compute $S^{\text{OBC}}$ which requires the determination of the non-universal constant $s_{1}$. Fortunately, $s_{1}$ can be easily extracted by comparing DMRG results with Eq. 8 for a periodic XXX chain. Based on this $s_{1}$ value, we find that, at the Kondo fixed point ($J_{K}=1$), $g\approx 1.623$ for $M=3$ and $g\approx 1.787$ for $M=4$. Both these values are very close to the predictions of Eq. 1, i.e., $g(M=3)=1.618$ and $g(M=4)=1.732$, even for the relatively small system sizes we used. The difference can be attributed to finite size effect as we will discuss below. The bond dimension $\chi$ needed is approximately $L^{\alpha\ c_{\text{eff}}}$ where $\alpha$ is a constant and $c_{\text{eff}}=M$ is the effective central charge. The computational cost is about $\mathcal{O}(L\times\chi^{3})\sim\mathcal{O}(L^{3\alpha M+1})$ [72], which scales exponentially with $M$.
@@
We seek to further support our proposed multichannel model by showing that it reproduces a similar scaling of $S_{\text{imp}}$ and $\xi_{K}$ (Kondo correlation length) [74] observed for the one- and two-channel models [32, 37]. We performed an analysis of $S_{\text{imp}}$ as $J_{K}$ was varied for different values of the system size $L$ to extract $\xi_{K}$ by fitting to the scaling form $S_{\text{imp}}(x/L,x/\xi_{K})$. The extracted $\xi_{K}$ values are then shown in the inset of Fig. 2(a) where our empirically measured $\xi_{K}$ indeed appears to scale as $e^{a/J_{K}}$. The resulting scaling law for $S_{\text{imp}}$ is shown in Fig. 2(a). In contrast to the Fermi-liquid power law $1/(x/\xi_{K})$ observed in the spin-chain 1CK model [59, 61, 32] and the non-Fermi liquid power law $\ln(x/\xi_{K})/(x/\xi_{K})$ for the spin-chain 2CK model [37], we find the non-Fermi liquid power law $1/(x/\xi_{K})^{4/5}$ in our spin-chain 3CK model, which is more dominant than the Fermi liquid $1/(x/\xi_{K})$ coming from each XXX chain. Both finite-size power laws match the finite-temperature correction to the multichannel impurity entropy [75, 76], which is related to the specific heat [16] when taking the temperature $T\to 1/(x/\xi_{K})$. Fig. 2(a) allows us to estimate the finite-size correction $\exp S_{\text{imp}}(x)-g(M)$. For $M=3$, $g(3)=\varphi$ and the parameters of Fig. 1(b), we can estimate this correction to be $\sim 0.01$ for $x=L/2$. Similar estimate for $M=4$ (albeit using the above $\xi_{K}$) gives $\sim 0.03\equiv\epsilon$, of the same order of magnitude as $S_{\text{imp}}-\ln(\sqrt{3})=\ln(\sqrt{3}+\epsilon)-\ln(\sqrt{3})\approx\epsilon/\sqrt{3}$ in Fig. 1(c). Similar to the above finite-size scaling is found for the total impurity spin, discussed next.
@@
V Total Impurity Spin